HUBBLE SPACE TELESCOPE DAILY REPORT #2189 PERIOD COVERED: 0000Z (UTC) 08/17/98 - 0000Z (UTC) 08/18/98 Daily Status Report as of 230/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/1/2 7258 (The Nature of the Compact Infrared Core Sources in Powerful FR II Radio Galaxies) The Near Infrared Camera (Camera One) was used to make observations of the compact infrared nuclear core region of the powerful radio galaxy 3C405. It is hoped that these observations will provide some of the most compelling evidence for the idea that all powerful radio galaxies contain luminous quasars which are hidden from our direct view by obscuring torii. The observations were completed as planned, and no anomalies were reported. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/2 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera Two) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.4 Completed NIC/1/2/3 7857 (Investigating the Missing Link Between Disks Around Pre-Main Sequence and Main Sequence Stars) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of star HD145788-C. Various observations and models of the circumstellar disk around the Main Sequence IR excess star Beta Pictors tend to show that the optically thin dust disk is the result of destruction processes among planetesimals. The observations were completed as planned, and no problems were reported. 1.5 Completed Seven Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. As described in HSTAR 6625 and 2.1, the acquisition for the fifth iteration of this proposal failed and the take data flag remained down for one observation. Otherwise, the observations were completed as planned, and no further problems were reported. 1.6 Completed NIC/2 7887 (A Survey of Gravitational Lenses as Cosmological Tools II) The Near Infrared Camera (Camera Two) was used to make observations of the gravitational lens candidate system HST154433+5. The propose is to extend our IR survey of gravitational lenses and gravitational lens candidates by obtaining NICMOS images of 14 additional lens systems or candidates. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Four Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. As described in HSTAR 6625 and 2.1, the acquisition for the third iteration of this proposal failed and the take data flag remained down for two observations. Otherwise, the observations were completed as planned, and no further problems were reported. 1.8 Completed STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.9 Completed Two Sets of NIC/2 7835 (A Search for Superplanets Embedded in Beta Pic & Vega-like Circumstellar Disks) The Near Infrared Camera (Camera Two) was used to make observations of the circumstellar disk around star HD10700 in order to search for superplanets. The discovery of a circumstellar disk orbiting Beta Pictoris as well as the discovery of other similar IR-excess Vega-like systems have provided compelling evidence that our solar system, or at least many of the processes which are thought to have created it, may be far from unique. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Three Sets of NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence star Alpha-PSA2. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.11 Completed FGS 7485 (Dynamical Masses for the Stars in the Pre-Main-Sequence Spectroscopic Binary 045251+3016) The Fine Guidance Sensors were to be used to make astrometric observations of the pre-main-sequence, spectroscopic, binary star STAR-04560 in order to obtain its dynamic mass. An excellent single-lined spectroscopic orbit is already completed, which provides all of the essential orbital elements except inclination angle and mass ratio. As described in HSTAR 6625 and 2.1, the acquisition for this proposal failed and the take data flag remained down. 1.12 Completed Two Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. As described in HSTAR 6625 and 2.1, the acquisition for the first iteration of this proposal failed and the take data flag remained down, affecting two observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.13 Completed Three Sets of STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The final iteration of this proposal was affected by the STIS suspend described in 2.3. Otherwise, the observations were completed as scheduled, and no further problems were reported. 1.14 Completed STIS/MA1 7645 (STIS FUV-MAMA Cycle 7 Flats) The Space Telescope Imaging Spectrograph (MA1) was to be used to make far-ultraviolet flat calibration images. This program will obtain FUV-MAMA flat-field observations with the Kr lamp for the construction of on-orbit D-flats for G140M. A dark calibration image was also to be taken. The observations were affected by the STIS suspend described in 2.3 which occurred near the beginning of this proposal. 1.15 Completed NIC/2 7895 (HST NICMOS SNAPSHOT Observations of Field Galaxies at Z < 1) The Near Infrared Camera (Camera Two) was used to make observations of the field galaxy GAL-221735. The proposed NICMOS observation will provide their morphology in rest frame NIR wavelengths {8000 - 16000 A}, where the effect of dust extinction is less severe, and their near infrared {NIR} colors. The observations were completed as scheduled, and no anomalies were reported. 1.16 Completed STIS/CCD/WF/PC-2/NIC/1/2/3 7174 (LBDS 53W091: A Cluster of Old, Red Galaxies) The Space Telescope Imaging Spectrograph (CCD), the WF/PC-2 and the Near Infrared Camera ( Cameras One, Two and Three) were used to make observations of the cluster of old red galaxies known as LBDS53W091. This distant cluster { at Z=1.55 } is representative of the oldest stellar systems we can glimpse at their "middle age"; and we are interested in their morphologies. The STIS portion if this proposal was affected by the STIS suspend described in 2.3. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 9 The acquisition at 229/180805Z failed when it was interrupted by astometry operations. A high attitude error vector was observed. HSTAR 6625 was written. The proposals described in 1.5, 1.7, 1.11, and 1.12 were affected. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 20 Successful : 20 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was reset twice. The PRS string (AP2) with the old PRS software was released at 229/1016Z. A DF-224 dump was collected at 229/1535Z on CCS strings A and B. As described in HSTAR 6626, a NICMOS single bit error occurred at 229/212356Z. At 229/223609Z the error log was dumped and the ground system limit was changed from three to four. Per HSTAR 6629, STIS suspended at 229/215515Z while at the beginning of the proposal described in 1.14. Indications are that the suspend was caused by an invalid MAMA optical mode selection during a MAMA1 high voltage operation. Also affected were the proposals described in 1.13 and 1.16, as well as any STIS proposal occurring before the STIS is recovered (which will be reported in tomorrow's report). Using ROP NS-10, the STIS memory was dumped at 229/2344Z, the SI C&DH errors were reset at 230/0126Z per ROP NS-5, and the NSSC-1 status buffer was dumped and reset at 230/0919Z per ROP NS-3. Recovery could be attempted at approximately 230/1049Z, although another occurrence of the erroneous commanding will occur at 230/1800Z. The anomaly and the future problem will be discussed at a meeting this morning. Real-time operations tranferred to the SAMS backup facility at 230/0600Z to verify the SAMS capability and to provide practical training for operations personnel. PASSOPS will remain in the MOR since UPS is not yet available in the SAMS facility. Operations personnel should return to the MOR at approximately 230/2400Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH